### v1

# A Survey of AI Scientists: Surveying the automatic Scientists and Research

Abstract
A new scientific paradigm, the AI Scientist, has coalesced at the intersection of artificial intelligence and epistemology, promising a fundamental shift from AI-assisted analysis to end-to-end autonomous discovery. Catalyzed by rapid advances in large language models, multi-agent orchestration, and robotic automation, these systems are architected to emulate the complete scientific workflow—from initial hypothesis generation to the final synthesis of publishable findings. This transition moves beyond using AI as an instrument of inquiry, positioning it as a potential originator of scientific knowledge. However, the rapid and unstructured proliferation of these systems has created a fragmented research landscape, obscuring overarching methodological principles and developmental trends. This survey provides a systematic and comprehensive synthesis of this emerging domain by introducing a unified, six-stage methodological framework that deconstructs the scientific process into: Literature Review, Idea Generation, Experimental Preparation, Experimental Execution, Scientific Writing, and Paper Generation. Through this analytical lens, we systematically map and analyze dozens of seminal works from 2022 to late 2025, revealing a clear three-phase evolutionary trajectory. Our analysis charts the field’s progression from an initial phase of Foundational Modules (2022–2023), focused on task-specific automation, through a period of Closed-Loop Integration (2024), to the current frontier of Scalability, Impact, and Collaboration (2025–present). By synthesizing these developments, this survey identifies key architectural patterns and highlights the dual research thrusts toward both greater machine autonomy and more sophisticated human-in-the-loop synergy. We conclude by presenting a forward-looking agenda that addresses critical open challenges in robustness, generalizability, and ethical governance. Ultimately, this work provides a critical roadmap for the field, intended to guide the next generation of systems toward becoming trustworthy, verifiable, and indispensable partners in human scientific inquiry. Project Github: https://github.com/Mr-Tieguigui/Survey-for-AI-Scientist.
33footnotetext: Latest Update: Oct., 2025.
Keywords AI Scientist, Autonomous Science, Large Language Models, Multi-Agent Systems, Scientific Workflow Automation.
1 Introduction
Over the past few years, a new paradigm has coalesced at the intersection of artificial intelligence and the philosophy of science: AI Scientist. Distinct from earlier "AI for Science" efforts that leveraged machine learning to accelerate discrete tasks like data analysis or simulation [1, 2, 3], the AI Scientist represents a transformative ambition toward end-to-end autonomous discovery. Rather than serving as computational assistants, these systems are architected to emulate and, in some cases, fully execute the roles of human researchers—formulating novel hypotheses, designing and conducting experiments, interpreting results, and generating publishable insights [4, 5, 6]. This vision has been catalyzed by a confluence of recent breakthroughs, including the sophisticated reasoning capabilities of large language models (LLMs) [7, 8], advances in multi-agent orchestration [9, 10], and the maturation of automated laboratory systems [11, 12]. Together, these developments are driving a fundamental transition from automation—where AI assists predefined steps—to autonomy—where an AI agent designs, validates, and executes its own scientific workflow. This conceptual shift not only reshapes how research is conducted but also poses profound epistemological questions: AI is evolving from an instrument of inquiry into a potential originator of scientific knowledge.
Historically, the idea of machine-driven science can be traced to early symbolic reasoning and automated theorem proving [13, 14]. Yet, it was constrained by domain specificity and limited generalization. Modern AI Scientists, powered by foundation models and reasoning frameworks, have broken this bottleneck. They integrate symbolic reasoning, natural language understanding, and multi-modal perception to create a self-improving research loop: observe–hypothesize–experiment–analyze–publish. Recent systems such as The AI Scientist [15], Curie [11], and PiFlow [16] exemplify this progression, combining agentic planning, principle-aware inference, and iterative self-correction. This evolution parallels milestones in autonomous experimentation—ranging from closed-loop chemical discovery [17] and bioinformatics workflows [18] to equation discovery [19]. The integration of these components points to an emerging research discipline, where scientific reasoning and empirical validation converge in fully autonomous frameworks.
Since 2024, the research community has seen an exponential rise in literature addressing AI-based scientific autonomy. Comprehensive surveys such as [4], [5], and [20] have highlighted this shift, while specialized frameworks such as The AI Scientist-v2 [6] and AI-Researcher [21] have demonstrated near-human performance in research ideation and experimental reasoning. Complementary efforts—e.g., Auto-Bench [22], ResearchBench [23], and IdeaBench [24]—have established benchmarks to measure the novelty, causality, executability, and reproducibility (NCER) of AI-driven discoveries. Despite this progress, existing studies remain fragmented. Most works either focus on specific domains (e.g., chemistry, biology, physics) or individual capabilities (e.g., hypothesis generation [25, 26], literature synthesis [27, 28]) rather than holistic frameworks. No prior survey has yet offered a unified taxonomy linking scientific tasks, AI capabilities, agentic systems, and evaluation protocols. This survey therefore aims to synthesize these disparate threads and establish a coherent foundation for the field.
This survey focuses on research between 2022 and 2025 that enables autonomous or near-autonomous execution of the scientific method. To ensure conceptual clarity and structural coherence, we adopt the radial organization illustrated in Figure 1, which conceptualizes the AI Scientist as a core engine driving a cyclical workflow that is instantiated across various scientific domains. At its heart lies the central concept of the Automatic AI Scientist System. Surrounding this core is an inner ring depicting the six fundamental methodological categories that constitute the scientific workflow: (1) Literature Review focuses on extracting, structuring, and reasoning over scientific corpora to establish prior knowledge foundations [27]; (2) Idea Generation addresses hypothesis formation and creative scientific reasoning [25, 26]; (3) Experimental Preparation encompasses data selection, variable definition, simulation setup, and statistical initialization [29, 30]; (4) Experimental Execution integrates protocol orchestration, robotic control, feedback-based iteration, and tool-use reasoning [12, 11, 31]; (5) Scientific Writing involves multi-modal evidence organization, figure and table generation, and result articulation [27, 32]; and (6) Paper Generation targets the synthesis of publishable, verifiable manuscripts that combine reasoning, visualization, and scholarly consistency [21, 33]. As shown in Figure 1, encircling this workflow is an outer ring representing the application layers, showcasing where these methodologies are applied—from general-purpose frameworks to specialized domains such as Chemistry, Biology, Physics, and Meta-Science. This concentric structure provides a unified, navigable map for understanding how a core autonomous system executes a fundamental scientific process across a diverse and expanding landscape of applications.
Figure 1: The concentric architecture of an AI Scientist System. The core represents the central Automatic AI Scientist System engine. This engine drives the six-stage methodological workflow, depicted in the inner ring, which forms a continuous, iterative cycle from Literature Review to Paper Generation. The outer ring illustrates the key application domains where this workflow is instantiated, from general-purpose frameworks to specialized fields like Chemistry, Biology, and Physics.
1.1 Contributions
This paper represents the first comprehensive survey on AI Scientists, providing a systematic, structured, and multidisciplinary synthesis of this rapidly emerging field. While earlier reviews have examined isolated aspects—such as LLM evaluation [9], hypothesis generation [25], or agentic automation frameworks [2]—none have yet captured the full landscape of autonomous scientific reasoning and experimentation. In contrast, this survey adopts an integrative perspective, combining methodological foundations, system-level architectures, and benchmark ecosystems within a unified framework, as illustrated in Figure 1. The main contributions are summarized as follows:
  • First Survey with Unified Six-Stage Framework. This paper presents the first comprehensive survey to introduce and systematically apply a principled, six-stage methodological framework that deconstructs the end-to-end scientific workflow. We model the process as a sequence of six interoperable stages: Literature Review, Idea Generation, Experimental Preparation, Experimental Execution, Scientific Writing, and Paper Generation. This taxonomy moves beyond ad-hoc descriptions of system capabilities to establish a unified conceptual vocabulary for the field. By formalizing the dependencies and information handoffs between stages, our framework provides a robust analytical lens to systematically categorize disparate systems, compare their architectural choices, and identify critical gaps in the pursuit of full-cycle scientific autonomy.
  • Comprehensive Synthesis of Systems. Furthermore, we provide a comprehensive synthesis of the systems and benchmarks that define the AI Scientist landscape. Our analysis systematically maps dozens of seminal works from 2022 to late 2025 onto our six-stage framework, culminating in a detailed panoramic matrix that visualizes the field’s capabilities and chronological evolution. This synthesis surveys cutting-edge applications, from general-purpose architectures like DeepScientist [34] to domain-specific breakthroughs in chemistry, biology, and physics, providing a deeply empirical grounding for the entire survey.
  • Identification of a Three-Phase Historical Trajectory. Based on our comprehensive analysis, we identify and articulate a clear three-phase historical evolution of the field: from Foundational Modules (2022-2023) to Closed-Loop Integration (2024), and finally to the current frontier of Scalability, Impact, and Collaboration (2025), exemplified by systems like DeepResearcher [35] and freephdlabor [36]. This narrative provides a coherent developmental arc for understanding the field’s past and future.
1.2 Organization
The remainder of this survey is structured to guide the reader logically through the AI Scientist landscape. Section 2 first establishes our six-stage taxonomy and presents a comprehensive matrix of key works, before outlining the field’s three-phase historical evolution. Building on this foundation, Section 3 provides a deep dive into the methodologies of each of the six stages. Section 4 surveys the practical application of these systems, covering both general-purpose architectures and domain-specific instances. Finally, Section 5 discusses open challenges and future research directions, and Section 6 concludes the survey.
2 Background and Taxonomy
This chapter establishes the conceptual framework for our survey by introducing a taxonomy of research categories and tracing the historical evolution of AI Scientist systems. We begin by deconstructing the end-to-end scientific process into six distinct methodological stages. This classification serves as the primary analytical lens through which we categorize existing works. We then present a comprehensive matrix aligning representative systems and benchmarks from 2022 to 2025 with these stages, built upon a rigorous, verifiable review of each cited paper. Finally, we synthesize these findings into a historical narrative, identifying three major phases of development that chart the field’s progression from modular tools to integrated, self-reflective research agents.
Table 1: Comprehensive matrix of AI Scientist works (updated to 2025) aligned with six methodological stages. This table has been rebuilt and expanded based on a rigorous, verifiable review of each paper’s primary contributions. Each colored symbol represents explicit coverage of a methodological stage.
Work Lit. Idea Exp. Exec. Writ. Paper Year
DS-1000 [37] 2023.04
Coscientist [38] 2023.06
BioPlanner [39] 2023.10
MLAgentBench [40] 2023.10
LitLLM [27] 2024.02
The AI Scientist (v1) [15] 2024.08
SciAgents [41] 2024.09
IdeaBench [42] 2024.11
Quantum-Agent-SDL [43] 2024.12
HypER [26] 2025.01
The AI Scientist (v2) [6] 2025.02
Curie [44] 2025.02
AI co-scientist [45] 2025.02
ResearchBench [23] 2025.03
DeepResearcher [35] 2025.04
AutoLabs [31] 2025.04
AI-Researcher [46] 2025.05
EXP-Bench [47] 2025.05
Agentic AutoSurvey [48] 2025.09
PiFlow [16] 2025.09
DeepScientist [34] 2025.09
SR-Scientist [19] 2025.10
Freephdlabor [36] 2025.10
2.1 Taxonomy of Research Categories
AI Scientist research from 2022 to 2025 can be systematically deconstructed into six methodological stages. Each stage represents a critical phase of the scientific process that has been progressively automated. Together, they form an end-to-end pipeline that transforms unstructured knowledge into verifiable scientific output, bridging abstract cognition with concrete execution and communication.
  • $\bullet$
    Literature Review (Lit.). This foundational stage involves transforming unstructured scientific corpora into machine-interpretable knowledge. It encompasses techniques from large-scale information retrieval to the synthesis of research gaps. Systems like LitLLM [27] and HypER [26] focus on citation-grounded summarization and knowledge extraction, while advanced frameworks like SciAgents [41] and DeepResearcher [35] utilize web-scale interactions and graph reasoning to map existing knowledge and provide a robust foundation for downstream tasks.
  • $\bullet$
    Idea Generation (Idea). Building upon the structured knowledge from the prior stage, this phase automates hypothesis discovery and problem formulation. It leverages the creative and reasoning capabilities of LLMs to propose novel yet plausible research directions. This capability is explicitly evaluated by benchmarks like IdeaBench [42], and is a core component of both domain-specific systems like Coscientist [38] and advanced end-to-end frameworks like DeepScientist [34].
  • $\bullet$
    Experimental Preparation (Exp.). This crucial intermediate stage translates an abstract hypothesis into an executable plan. It includes tasks such as defining variables, selecting datasets, generating analysis code, and designing experimental protocols. This is a primary focus of data-science-oriented benchmarks like DS-1000 [37] and MLAgentBench [40], and is a key step in all integrated systems from The AI Scientist v1 [15] to freephdlabor [36].
  • $\bullet$
    Experimental Execution (Exec.). This stage involves the actual running of real or simulated experiments. It emphasizes the agent’s ability to interact with tools, control robotics, and adapt its plan based on real-time feedback. Milestones in this area include systems that orchestrate physical laboratory hardware, such as Coscientist [38] and Quantum-Agent-SDL [43]. Frameworks like Curie [44] and DeepResearcher [35] demonstrate this capability in simulated and real-world web environments, respectively.
  • $\bullet$
    Scientific Writing (Writ.). This stage focuses on the communication of scientific findings by transforming structured results into coherent, citation-grounded narratives. Capabilities range from section-aware summarization to data-to-text synthesis. This is a key feature in end-to-end systems like ResearchBench [23] and is central to human-in-the-loop frameworks like freephdlabor [36], where the AI drafts content for human review and refinement.
  • $\bullet$
    Paper Generation (Paper). Representing the culmination of the scientific workflow, this final stage synthesizes a full, publication-ready manuscript. This requires the tight integration of all prior stages. This end-to-end capability is the hallmark of the most advanced, fully autonomous systems, such as The AI Scientist v1/v2 [15, 6], AI-Researcher [46], and DeepScientist [34].
Figure 2: Evolution of AI Scientist research (2022–2025). A horizontal timeline illustrates three major phases: (I) Foundational Modules (2022–2023), (II) Closed-Loop Integration (2024), and (III) The Frontier: Scalability, Impact, and Collaboration (2025–present). Each phase highlights representative systems, with upward arrows denoting increasing levels of autonomy and integration.
2.2 Historical Evolution
The evolution of AI Scientist research from 2022 to 2025 reveals a clear trajectory: a progressive integration of automation, moving from discrete, task-specific modules toward self-reflective and verifiable end-to-end systems. We identify three major developmental phases that capture this conceptual deepening.
Phase I: Foundational Modules (2022–2023).
This initial phase was characterized by the development and benchmarking of components that address specific stages of the scientific process. Early works like DS-1000 [37] focused on the Experimental Preparation and Execution stages for data science code. In parallel, systems like BioPlanner [39] tackled protocol planning, while Coscientist [38] demonstrated the feasibility of closed-loop Execution in a physical lab. Benchmarks such as MLAgentBench [40] began to formalize the evaluation of these modular capabilities.
Phase II: Closed-Loop Integration (2024).
The year 2024 marked a critical turning point where the field shifted its focus from modular components to the integration of multiple stages into continuous workflows. A proliferation of specialized systems for cognitive tasks emerged, such as SciAgents [41] for Idea Generation. The milestone of this phase was The AI Scientist v1 [15], which successfully demonstrated the first fully autonomous, end-to-end Paper Generation loop, unifying the previously disparate stages into a single, cohesive process.
Phase III: The Frontier: Scalability, Impact, and Collaboration (2025–present).
The most recent phase is defined by three distinct and parallel research thrusts at the frontier of autonomous science. The first is the pursuit of scalability and robustness through deep learning. DeepResearcher [35] epitomizes this by using reinforcement learning in real-world web environments to train agents that can handle noisy, unstructured information, thereby improving over time. The second thrust targets scientific impact and progressive discovery. DeepScientist [34] is a landmark system designed for goal-oriented, long-horizon research with the explicit aim of surpassing the human state-of-the-art on frontier scientific tasks. The third, and equally significant, trend is towards deep human-AI collaboration. Frameworks like freephdlabor [36] architect the research process as a continual and interactive partnership, where a human researcher can guide, customize, and collaborate with a personalized multi-agent team. This evolution towards more scalable, impactful, and collaborative systems marks the growing maturity of the field.
Figure 3: End-to-end workflow of an AI Scientist system. The six stages represent a closed scientific loop, starting from knowledge synthesis and ending with validated scientific reports. Arrows denote data and reasoning flow, while the outer frame indicates embedded reflection and evaluation mechanisms.
3 Methodological Integration of AI Scientist Systems
This chapter provides a systematic analysis of the methodological components that constitute modern AI Scientist systems. Building upon the taxonomy in Table 1, we organize the end-to-end research workflow into six sequential stages: Literature Review (Sec 3.1), Idea Generation (Sec 3.2), Experimental Preparation (Sec 3.3), Experimental Execution (Sec 3.4), Scientific Writing (Sec 3.5), and Paper Generation (Sec 3.6). Each stage corresponds to a specific part of the scientific process, collectively forming a closed-loop pipeline for autonomous scientific discovery, as illustrated in Figure 3.
3.1 Literature Review
The Literature Review stage constitutes the foundational cognitive layer of an AI Scientist system. Its primary objective is to transmute vast, unstructured scientific corpora—including papers, protocols, and databases—into structured, provenance-aware knowledge representations. These representations are engineered to support high-level downstream tasks such as logical reasoning and hypothesis formation. Diverging from generic retrieval or summarization, an automated scientific literature review imposes stringent requirements for scientific faithfulness, verifiable citation grounding, and knowledge-level abstraction. This necessitates a system capable of reasoning over the established prior art, rather than merely condensing it. Consequently, this stage demands the sophisticated integration of Information Retrieval (IR), Natural Language Processing (NLP), and symbolic reasoning techniques into a unified pipeline that guarantees completeness, interpretability, and auditable provenance. To ensure methodological rigor, reproducibility, and modular design, the literature review process can be formalized as the five-stage pipeline depicted in Figure 4. Each consecutive stage constructs a progressively higher order of semantic structure upon the same textual substrate, addressing distinct computational objectives, data representations, and algorithmic challenges.
Figure 4: Pipeline for Automated Literature Review. The workflow begins with large-scale corpus ingestion (P1) and progresses through retrieval & re-ranking (P2), evidence structuring (P3), schema induction (P4), and grounded narrative synthesis (P5). Dashed arrows represent feedback loops: knowledge-gap identification in P5 can trigger targeted evidence retrieval in P2 or structure refinement in P3. The modular design supports both batch and agentic implementations.
  • $\bullet$
    Stage 1: Corpus Acquisition and Layout-Aware Parsing. The initial stage focuses on constructing a structured and queryable substrate from heterogeneous scientific documents. The process commences with the ingestion of large-scale document repositories (e.g., arXiv, PubMed, Semantic Scholar), followed by a normalization phase to standardize metadata, sectional hierarchies, and citation formats. Core methodologies include: (1) layout-aware parsing, which employs visual segmentation and optical character recognition (OCR) to reconstruct the logical reading order and preserve structural semantics; (2) conversion of raw PDF documents into standardized formats like TEI/XML or JSONL, facilitating granular indexing of sections, figures, or equations; and (3) reconstruction of the citation graph, linking inline references to bibliographic entries to enable robust provenance tracking. Foundational toolchains such as S2ORC [49] and GROBID [50] have been instrumental in this domain, providing large-scale, full-text corpora enriched with precise entity and citation annotations. The output is a canonical, layout-aware text collection that maintains the intrinsic spatial and relational architecture of scientific publications.
  • $\bullet$
    Stage 2: Retrieval and Re-Ranking. With a structured corpus in place, the system proceeds to retrieve and rank textual passages pertinent to a specified research question or hypothesis. This critical step must carefully balance recall and precision through the integration of multiple retrieval paradigms. The typical workflow involves: (1) Hybrid retrieval, which synergizes sparse lexical methods (e.g., BM25) with dense semantic encoders (e.g., SciBERT, SPECTER2) to capture both terminological specificity and conceptual relevance. (2) Cross-encoder re-ranking, which subsequently refines the candidate pool using contextual attention over query-passage pairs, often fine-tuned on scientific corpora like SciDocs to enhance domain coherence. (3) Section-aware weighting, which introduces discourse-level priors, prioritizing passages from high-signal sections such as abstracts or results. Contemporary systems like PaperQA [51] and its successor, which demonstrated superhuman synthesis capabilities [52], advance this stage by integrating Retrieval-Augmented Generation (RAG). This technique injects top-ranked passages directly into Large Language Model (LLM) prompts to facilitate citation-grounded synthesis. The final output is a ranked list of evidence snippets, each annotated with relevance and confidence scores for subsequent processing.
  • $\bullet$
    Stage 3: Evidence Structuring and Representation Learning. This stage orchestrates the transformation of retrieved textual evidence into machine-interpretable structures, bridging the gap between unstructured narrative and formalized scientific knowledge. The conversion is typically realized through three complementary approaches: (1) Information Extraction (IE), which identifies and normalizes key entities, relations, and measurements, yielding atomic tuples that encapsulate experimental claims. (2) Knowledge Graph (KG) induction, which aggregates these tuples into domain-specific graphical models—linking hypotheses, methods, and outcomes—and grounds them in established ontologies like UMLS or MeSH. (3) Table synthesis, which aligns disparate statements into comparative matrices, standardizing variables and metrics to support systematic reasoning. For instance, the Text-Tuple-Table (T3) framework [53] formalizes a text-to-tuple-to-table mapping using constrained decoding, while TKGT [54] demonstrates that integrating textual evidence with intermediate graph construction enhances factual coherence. This process yields a set of structured representations—tuples, graphs, or tables—each explicitly linked to its source provenance within the corpus.
  • $\bullet$
    Stage 4: Schema Induction and Comparative Table Construction. Following evidence structuring, the system induces abstract schema templates to identify meaningful dimensions for comparison across multiple research papers. The core challenge lies in the automated discovery of salient attributes—such as "model architecture," "dataset," or "evaluation metric"—that should form the columns of a comprehensive review table. This is operationalized through a multi-step process: (1) aspect clustering, which groups semantically similar textual mentions using sentence embeddings or contrastive encoders; (2) column induction, often performed via LLM prompting conditioned on topic distributions to propose a consistent set of table headers; and (3) cell population, which utilizes retrieval-conditioned generation to populate the schema slots with extracted facts or numerical data. The ArxivDIGESTables [55] system provides a complete implementation of this pipeline, effectively decomposing schema learning and value filling into two distinct supervised subtasks. The resulting output is a comparative survey table that aligns multiple studies along unified methodological axes, creating a structured knowledge base for both human and agentic consumption.
  • $\bullet$
    Stage 5: Grounded Narrative Synthesis and Novelty Analysis. The final stage synthesizes the structured evidence into a coherent, citation-grounded narrative, while concurrently quantifying the novelty of findings relative to prior art. This process integrates the structured graphs and tables back into textual reasoning via several mechanisms: (1) retrieval-conditioned generation, where every generated claim is constrained by and anchored to explicit evidentiary sources and citations; (2) contrastive novelty modeling, which compares generated summaries against literature-derived knowledge graphs to highlight unexplored parameter spaces or identify missing relational links; and (3) factuality verification, implemented through iterative retrieval-generation feedback loops designed to detect hallucinations and enforce comprehensive citation coverage. Systems such as LitLLM [56] implement robust pipelines for citation-grounded drafting, whereas SCIMON [57] specifically focuses on optimizing for inspiration diversity and novelty through iterative refinement. The outputs of this stage—structured narratives and knowledge-gap maps—explicitly articulate what is known, what remains uncertain, and where scientific opportunities reside, thereby setting the stage for hypothesis generation.
3.2 Idea Generation
Following the synthesis of existing literature, the Idea Generation phase functions as the creative nexus of the AI Scientist pipeline, as illustrated in Figure 5. Its purpose is to transform the structured knowledge, graphical representations, and literature embeddings produced in the prior stage into concrete, testable hypotheses and novel research directions. This task transcends generic text generation or summarization, demanding a synthesis of scientific creativity, semantic grounding, and novelty control. The system must not only forge unseen conceptual connections but also maintain rigorous epistemic validity. Recent research has formalized this process into a sequence of distinct reasoning and generation stages, encompassing conceptual fusion, cross-domain extrapolation, multi-agent brainstorming, and hypothesis scoring [58, 59]. This section delineates a four-stage methodological framework that integrates knowledge-driven reasoning, multi-agent collaboration, and evaluative feedback to facilitate autonomous hypothesis discovery.
Figure 5: Pipeline for Idea Generation. The process flows from conceptual fusion and trend extrapolation, through hypothesis refinement, multi-agent brainstorming, and final scoring & prioritization of hypotheses.
  • $\bullet$
    Stage 1: Conceptual Fusion and Trend Extrapolation. This initial stage is dedicated to uncovering latent connections within and across heterogeneous scientific corpora and identifying emergent research trajectories. Methodologies often employ dense retrieval and embedding-based clustering to map conceptual relationships between distinct research fronts [58, 60]. More advanced, domain-specific systems such as MOOSE-Chem utilize chemical knowledge graphs to rediscover unseen molecular relations, thereby demonstrating the capacity of LLMs for creativity within constrained scientific domains [61]. Concurrently, temporal citation analysis and trend forecasting techniques, such as dynamic topic modeling and citation-burst detection, guide the extrapolation process toward under-explored problem areas. By systematically constructing a large-scale "hypothesis bank"—annotated with novelty metrics and provenance scores—this stage establishes a robust foundation for subsequent refinement.
  • $\bullet$
    Stage 2: Hypothesis Refinement and Knowledge-Grounded Pruning. Once a corpus of initial hypotheses has been generated, this stage applies rigorous refinement and pruning to ensure their logical coherence, factual consistency, and empirical feasibility. Knowledge-grounded approaches combine LLM reasoning with external knowledge graphs to perform validation against established facts and causal principles [62]. Systems like HypER introduce provenance-aware distillation, which explicitly traces each hypothesis back to its source literature to enhance verifiability [63]. In specialized domains such as biomedicine, the integration of simulation feedback loops has proven effective for refining biological hypotheses [64]. The refinement process typically employs a combination of techniques, including: (1) reinforcement learning with iterative prompting to balance novelty and validity; (2) counterfactual simulation to stress-test causal chains; and (3) uncertainty-aware rejection sampling to filter out non-viable or poorly-supported propositions.
  • $\bullet$
    Stage 3: Multi-Agent Brainstorming and Cross-Domain Recombination. Scientific innovation often arises from the dialectical process of expert dialogue. AI Scientist systems emulate this dynamic by orchestrating collaborative LLM agents. Multi-role architectures—assigning agents to functions such as critic, generator, and verifier—have been shown to yield more diverse and impactful ideas than monolithic agent systems [59, 65]. Frameworks like Scideator operationalize human-LLM co-ideation by recombining conceptual facets across paper structures [66], while Nova applies iterative planning to systematically enhance the novelty and diversity of generated ideas [67]. These agents typically communicate through structured message passing grounded in scientific knowledge graphs [68], creating a networked reasoning ecosystem that promotes divergent exploration while ensuring convergent synthesis.
  • $\bullet$
    Stage 4: Feasibility and Novelty Evaluation. The final stage is devoted to the systematic evaluation and prioritization of candidate hypotheses based on their scientific merit, originality, and potential for experimental execution. Lightweight, graph-based evaluation frameworks such as GraphEval have been developed to combine semantic-distance metrics with graph-coverage analysis for efficient assessment [69]. Concurrently, a suite of dedicated benchmarks—including IdeaBench [42], AI Idea Bench 2025 [70], and LiveIdeaBench [71]—has emerged to quantify human-aligned novelty and plausibility via large-scale expert annotation. Other benchmarks like ResearchBench introduce task-decomposition paradigms to evaluate inspiration-based idea generation [72]. Collectively, these methods provide the quantitative and qualitative signals necessary for ranking and selecting high-value hypotheses, thereby enabling a closed-loop transition to the experimental preparation phase.
3.3 Experimental Preparation
The Experimental Preparation module constitutes the pivotal transitional phase that bridges abstract hypothesis and empirical validation, transforming conceptual proposals into executable and auditable experiments. Within a general-purpose AI Scientist architecture, this stage serves as the operational backbone for translating theoretical insights into robust testing pipelines. It unifies critical sub-tasks, including dataset selection, environment configuration, execution control, and reproducibility management. As depicted in Figure 6, we formalize this process as a domain-agnostic, four-stage workflow applicable across diverse scientific fields such as physics, chemistry, biology, materials science, and data science. These stages—Experimental Framing, Environment & Instrumentation Setup, Protocol Implementation, and Reproducibility & Lifecycle Tracking—collectively reconcile the conceptual with the empirical, ensuring that the path to discovery is governed by scientific integrity and interpretability.
Figure 6: Pipeline for Experimental Preparation. A four-stage, domain-agnostic process that transforms abstract hypotheses into reproducible experiments: (1) Framing formalizes variables and measurable goals using structured and visual reasoning; (2) Environment & Instrumentation setups integrate data backends and visualization pipelines; (3) Protocol Execution operationalizes controlled, adaptive experimentation; and (4) Reproducibility ensures provenance, robustness, and cross-agent comparability.
  • $\bullet$
    Stage 1: Experimental Framing. This initial stage involves the translation of a conceptual hypothesis into an actionable and precisely specified experimental plan. The AI Scientist is tasked with formalizing key components such as independent and dependent variables, measurement protocols, evaluation metrics, and underlying theoretical assumptions, while rigorously grounding them in structured data representations. Recent work underscores that effective framing hinges on the ability to reason over heterogeneous structured information, including tables, charts, and scientific metadata. Benchmarks like TableBench [73] and frameworks such as Chain-of-Table [74] have shown that the adaptive transformation of tabular data enhances analytical coherence. Similarly, visual analytics systems like ChartQA [75] and ChartX [76] illustrate how multimodal comprehension—synergizing numerical and visual reasoning—improves experimental design by accurately identifying relevant variables and their inter-dependencies. In data-centric sciences, the automated translation of hypotheses into model templates, as explored in DiscoveryBench [77] and by Li et al. [78], demonstrates how LLMs can map theoretical relations to measurable quantities under formal statistical constraints. Collectively, these methodologies affirm that experimental framing can be conceptualized as an optimization problem: the selection of controllable variables and observable targets to maximize expected information gain while balancing constraints of cost and uncertainty.
  • $\bullet$
    Stage 2: Environment and Instrumentation Setup. Subsequent to the framing stage, the system establishes a controlled experimental environment—be it physical, simulated, or hybrid—engineered to ensure the fidelity and traceability of outcomes. The increasing integration of agentic pipelines for data analysis, benchmarked by suites like DSBench [79], BLADE [80], and InfiAgent-DABench [81], provides compelling evidence that modern scientific automation requires support for realistic computational backends where agents interact with diverse tools, databases, and sensors. Relational database-driven systems, exemplified by DAgent [82], show that structured schema reasoning and SQL-bound report generation can function as reliable surrogates for physical laboratory data flows, enabling autonomous experiment monitoring. Furthermore, visualization-aware toolchains such as AutomaTikZ [83] and Text2Chart31 [84] indicate that the pre-registration of analytical visual outputs (e.g., charts, plots, schematics) prior to execution enhances both interpretability and pre-execution validation. This stage culminates in a unified orchestration layer that consolidates instrumentation calibration, environment sandboxing, and visualization configuration, allowing the AI Scientist to deploy a reproducible experimental backbone.
  • $\bullet$
    Stage 3: Protocol Implementation and Plan Reasoning. This phase concerns the operationalization of the experiment, translating the abstract specification into a step-wise, executable routine designed for autonomous execution with adaptive feedback control. An emerging line of research in modeling-driven reasoning, epitomized by MM-Agent [85], demonstrates that formal mathematical modeling can serve as a universal backbone across scientific domains, translating conceptual relationships into solvable optimization or simulation tasks. Empirical evaluation frameworks, including DiscoveryBench [77] and DS-Agent [30], further affirm that multi-step execution plans guided by intermediate evaluation checkpoints consistently outperform single-pass strategies in both reliability and interpretability. At a systems level, realistic data-science benchmarks such as DSBench [79] and BLADE [80] highlight the imperative of handling heterogeneous data streams and dynamic schema evolution during runtime. These findings collectively suggest that robust protocol execution necessitates a capacity for continuous introspection, wherein an AI Scientist must not only perform prescribed tasks but also dynamically assess intermediate consistency, adapt hyperparameters, and re-execute failed components within a persistent experimental context.
  • $\bullet$
    Stage 4: Reproducibility and Lifecycle Tracking. The cornerstone of scientific credibility lies in the ability to reproduce and audit findings across different agents and timeframes. In this stage, the AI Scientist is responsible for maintaining detailed provenance, tracking parameter modifications, and enforcing consistent experimental states throughout the research lifecycle. Studies dedicated to benchmarking agentic systems reveal that transparency and provenance tracing are integral to reproducible evaluation; for example, BLADE [80] mandates detailed logging of environmental context and tool interactions, while DSBench [79] quantifies reproducibility deficits via cross-agent consistency scores. Complementary research in multimodal evaluation, exemplified by CharXiv [86] and the work of Deng et al. [87], has shown that performance variance across modalities can expose latent instabilities, thereby offering a blueprint for robustness testing in complex scientific workflows. Through automated provenance documentation, environment snapshots, and periodic re-validation, this stage ensures that the AI Scientist transitions from a heuristic experimenter to a verifiable scientific operator, whose outputs are primed for rigorous execution and analysis.
3.4 Experimental Execution
The Experimental Execution stage represents the empirical core of an AI Scientist system, where the previously designed protocol is operationalized. This phase translates abstract experimental plans into tangible actions, whether in real or simulated environments, to yield verifiable outcomes. In contrast to traditional, linear automation pipelines, this stage is characterized by its emphasis on closed-loop reasoning, multimodal monitoring, and dynamic self-correction mechanisms that continually link empirical outcomes back to the system’s conceptual model [38, 44, 47, 88]. As Figure 7 illustrates, this process can be deconstructed into a canonical four-stage pipeline: (1) Protocol Instantiation, (2) Instrument and Tool Invocation, (3) Adaptive Execution and Feedback, and (4) Data Acquisition and Validation. Together, these components bridge the divide between symbolic reasoning and tangible experimentation.
Figure 7: Pipeline for Experimental Execution. A unified four-stage process for AI Scientist systems: (1) Protocol Instantiation; (2) Instrument and Tool Invocation; (3) Adaptive Execution and Feedback; (4) Data Acquisition and Validation. Arrows indicate information flow and feedback loops between stages; blue paths denote automated data streams, and orange paths denote agentic control signals.
  • $\bullet$
    Stage 1: Protocol Instantiation. At the inception of the execution phase, the AI Scientist translates an abstract experimental design into a set of executable, domain-grounded procedures. In the biochemical domain, BioPlanner demonstrates how language models can convert natural-language plans into structured protocols suitable for bench-top validation [39]. In the context of causal discovery, research by Li et al. shows that LLMs can formulate and parameterize interventional study designs, thereby bridging high-level hypothesis formation with concrete experimental realization [89]. Systems like Curie formalize this entire process into modular representations that facilitate task decomposition, ensuring both traceability and reproducibility [44]. Recent advancements in hierarchical representations for protocol design further offer methods to encapsulate procedural templates, safety constraints, and device mappings, creating a structured abstraction layer that precedes physical execution [90].
  • $\bullet$
    Stage 2: Instrument and Tool Invocation. Once the protocol is instantiated, the AI Scientist must interface with a heterogeneous array of toolchains—including robotic laboratories, high-fidelity simulators, or computational clusters—to execute the prescribed tasks. The burgeoning field of self-driving laboratories (SDLs) provides foundational architectures for this stage, integrating robotic manipulation, spectroscopy, and closed-loop feedback control [91, 92]. Systems such as ORGANA [93] and AutoLabs [31] exemplify the multi-agent orchestration of robotic chemistry, where individual agents are delegated responsibilities like apparatus scheduling and calibration. In high-energy physics, agentic models have been shown to orchestrate multi-stage experiments at large-scale accelerator facilities, establishing the feasibility of autonomous control over complex physical instrumentation [94]. This stage necessitates the seamless integration of API calls, sensor feedback loops, and parallel instrument execution, unifying disparate control layers through structured metadata.
  • $\bullet$
    Stage 3: Adaptive Execution and Feedback. A core tenet of modern experimental execution is its departure from linear, pre-scripted workflows; instead, it demands real-time assessment and dynamic revision. Dynamic execution frameworks, such as the EXP-Bench benchmark [47] and the self-corrective control loops within Curie [44], demonstrate how intermediate evaluation metrics (e.g., error bounds, data quality, reproducibility scores) can be used to automatically adjust parameter ranges or trigger corrective strategies. This adaptive behavior is simulated in data-science contexts through looped code-generation-testing cycles, as demonstrated in benchmarks like DS-1000 [37] and by Yin et al. [95]. In chemical synthesis, the CoScientist system embodies this philosophy by continuously refining reagent ratios based on prior outcomes, effectively realizing a closed-loop reasoning process in the physical world [38]. These models illustrate how reinforcement, self-critique, and error detection converge to form a coherent adaptive control system.
  • $\bullet$
    Stage 4: Data Acquisition and Validation. This concluding stage focuses on the consolidation of raw experimental outputs into validated, provenance-linked datasets. The process involves multi-modal data capture (e.g., text, images, signals), comprehensive metadata annotation, and rigorous validation against both statistical and symbolic criteria [88]. For example, both Curie [44] and EXP-Bench [47] embed runtime analytics modules that validate experimental outputs against ground-truth simulations or established baseline models. In physics, this approach has been extended to large-scale sensor networks, where aggregated instrument telemetry is used for post-hoc causal interpretation [94]. Ultimately, robust data validation is the critical step that enables the subsequent stages of analysis and scientific writing, creating an unbroken chain of evidence by linking execution logs with the cognitive provenance trails established across the entire scientific pipeline.
3.5 Scientific Writing
The Scientific Writing stage constitutes the communicative endpoint of an AI Scientist system, where structured analytical findings are transformed into coherent, verifiable, and ethically compliant scholarly narratives. Diverging from generic text generation, this phase must ensure non-negotiable standards of factual grounding, data–text alignment, and publication-grade integrity. Recent studies confirm that large-language-model (LLM)–assisted writing has permeated nearly all domains of research communication, from biomedical publishing [96, 97] to educational and simulation science [98, 99]. The resulting ecosystem, summarized in Figure 8, integrates five interlinked components—drafting, data–text linking, peer-review automation, ethical governance, and publication optimization—which collectively form a closed feedback loop between human authors and their AI assistants.
Figure 8: Pipeline for Scientific Writing. The pipeline consists of five interlinked stages. Forward arrows (blue) represent the content generation flow, while feedback loops (orange/green) indicate review and governance feedback returning to the drafting stage. The visualization maintains the same visual style as previous subsections, using a minimalist blue–gray tone and rounded modular blocks.
  • $\bullet$
    Stage 1: Drafting and Structural Composition. At its foundation, the scientific writing process involves the conversion of analytical outputs into well-structured academic prose that adheres to established disciplinary conventions. LLM-based drafting systems, as surveyed in foundational works [100, 101], demonstrate significant capabilities in logical structuring, linguistic clarification, and maintaining cross-sectional coherence. In specialized fields, domain-specific assistants have been shown to achieve fluency comparable to expert authors when guided by section-specific prompts and retrieval-grounded content alignment [99, 96]. To mitigate hallucination and ensure factual reliability, citation-aware retrieval mechanisms map claims to verified sources, while structure-aware templates enforce parallelism between sections such as "Methods" and "Results" [102, 103]. Empirical studies further indicate that human-AI co-editing workflows can reduce grammatical noise and accelerate drafting efficiency [104].
  • $\bullet$
    Stage 2: Data–Text Linking and Multimodal Representation. Modern scientific communication is increasingly contingent upon the robust alignment between quantitative data and its descriptive narrative. Contemporary AI writing systems integrate sophisticated data-to-text pipelines that can automatically cross-reference tables, figures, and statistical outputs with their corresponding textual segments [102, 32]. Hybrid language-vision models are now capable of generating or refining visual elements—such as plots, charts, or TikZ diagrams—from structured data, thereby enhancing semantic coherence across modalities [105]. Furthermore, grounded captioning and multimodal feedback mechanisms enable the automatic rewriting of captions, suggestion of figures, and summarization of tables, ensuring that each visual element is presented with accurate and sufficient context [33]. These advances are redefining scientific writing as a form of multimodal composition, wherein LLMs act as mediators between quantitative data, visual analytics, and textual exposition.
  • $\bullet$
    Stage 3: Peer-Review Automation and Quality Verification. A significant frontier in AI-augmented scholarship is the automation of peer review. LLM-based reviewers, trained on extensive corpora of editorial decisions, can identify inconsistent claims, missing citations, and logical fallacies within a manuscript [106, 33]. Experimental systems now emulate human referee workflows by providing structured, actionable comments, quantitative quality scores, and explainable rationales for their assessments. The implementation of iterative "review–respond–revise" loops has the potential to significantly reduce latency in academic publishing while maintaining quality parity with human referees [97, 107]. Nevertheless, issues of reproducibility and algorithmic bias remain open challenges, prompting calls for transparent confidence calibration and greater reviewer accountability [100].
  • $\bullet$
    Stage 4: Ethical Compliance and Authorship Governance. Ethical oversight has become an integral component of the AI-assisted publication pipeline. Emerging policies consistently emphasize the mandatory disclosure of AI involvement and the explicit delineation of authorship to prevent ghostwriting and misrepresentation [108, 109, 110]. Educational and professional bodies are actively developing standardized checklists and guidelines—such as the Transparent Disclosure Protocol—for systematically documenting prompt histories, model identifiers, and revision provenance [111, 112]. Concurrently, scholars advocate for layered authorship attribution models that formally recognize both human intellectual contribution and algorithmic assistance [113, 103]. These governance mechanisms are essential for anchoring trust, accountability, and reproducibility in the evolving landscape of AI-augmented authorship.
  • $\bullet$
    Stage 5: End-to-End Publication Optimization. The final stage integrates the preceding components—drafting, multimodal generation, review, and ethics—into a seamless, AI-assisted publication workflow. Comprehensive frameworks, exemplified by the benchmarks WritingBench [32] and SPOT [33], are designed to evaluate manuscripts holistically for factual accuracy, citation soundness, and stylistic alignment with target journal standards. In such systems, adaptive agents monitor the document lifecycle, suggesting figure–text adjustments, formatting references, and validating ethical disclosures. By embedding these optimization modules within authoring environments, these systems empower scientists to transition from manual composition to a role of supervisory orchestration. This shift redefines scientific publication as a co-creative, auditable, and integrity-preserving process [97, 101].
3.6 Paper Generation
The Paper Generation stage represents the apex of an AI Scientist system’s capabilities, signifying a transition from assisting human researchers to autonomously crafting complete scientific manuscripts. This culminating phase integrates all preceding functionalities—ideation, experimentation, data visualization, text composition, and review simulation—into a single, unified workflow, as depicted in Figure 9. Exemplary systems such as The AI Scientist v1 and v2 have demonstrated this capacity by generating entire scientific papers, inclusive of figures, experiments, and internal reviews, with minimal human oversight [15, 6]. Emerging frameworks like AI-Researcher are further extending this paradigm by developing end-to-end research pipelines and associated benchmarks (e.g., Scientist-Bench) to systematically evaluate autonomous science workflows [46]. Herein, we decompose the paper-generation pipeline into four sequential stages: (1) Manuscript Drafting; (2) Visual & Tabular Composition; (3) Review & Revision Agent; and (4) Publication Dissemination.
Figure 9: Pipeline for Paper Generation. The workflow spans four sequential stages: Manuscript Drafting, Visual & Tabular Composition, Review & Revision Agent, and Publication Dissemination. Solid blue arrows represent the forward generation flow and thinner orange feedback arrows indicate revision loops between the visual/tablular and review stages.
  • $\bullet$
    Stage 1: Manuscript Drafting. In this initial phase, the system transforms structured analytical outputs—such as hypotheses, results tables, and code logs—into the full textual sections of a scientific manuscript. The AI Scientist v1, for instance, converts raw experiment logs into LaTeX-formatted drafts, retrieves relevant literature to synthesize context, composes "Related Work," "Methods," "Results," and "Discussion" sections, and automatically embeds the corresponding citations [15]. Subsequent research, such as the AIGS (AI-Generated Science) project, has explored automated falsification and multi-agent pipelines to achieve full-cycle manuscript generation beyond narrow domains [114]. Key enabling technologies at this stage include retrieval-augmented generation for contextually rich sections, section-aware language modeling to ensure inter-section coherence, and template-based formatting to adhere to specific journal or conference styles.
  • $\bullet$
    Stage 2: Visual & Tabular Composition. Once the textual draft is prepared, the system must generate or integrate essential visual artifacts—figures, charts, and tables—and ensure their seamless alignment with the narrative. In the workflow of The AI Scientist, experimental outputs are programmatically converted into plots and tables, for which captions are generated and cross-references are inserted directly into the LaTeX manuscript [15]. Multi-agent systems like SciSciGPT have demonstrated how to integrate data processing and manuscript drafting, including the automated generation of figures and the establishment of narrative links across different modalities [115]. Core technologies include automatic plotting from data or code outputs, figure-caption co-generation, dynamic table population, and robust LaTeX integration, all underpinned by multimodal consistency checks between text, tables, and figures.
  • $\bullet$
    Stage 3: Review & Revision Agent. Recognizing that scientific publication is contingent upon peer review, advanced systems embed internal reviewer agents to critique drafts, propose revisions, and iteratively refine the manuscript. For example, The AI Scientist v2 implements an agentic tree-search architecture that incorporates a dedicated reviewer-agent component, which is capable of evaluating drafts and generating structured feedback before the manuscript is finalized [6]. Similarly, frameworks such as AI-Researcher include modules designed for automated revision loops and the benchmarking of autonomous review generation [46]. Key methodologies in this stage include models trained on historical peer-review data, LLM-driven comment generation, automated revision implementation (for both text and figures), and iterative loop control that simulates author-reviewer interactions.
  • $\bullet$
    Stage 4: Publication Dissemination. In the final stage, the system prepares the fully revised manuscript for submission. This involves formatting the document according to target venue templates, managing all necessary metadata (e.g., authors, affiliations, disclosures), verifying ethical compliance, and potentially interfacing with submission systems via APIs. Industrial research efforts, such as Google’s work on empirical-software systems, are exploring automated code-to-paper pipelines that ensure submission readiness across various domains [116]. This pipeline also encompasses post-publication activities, such as versioning, citation tracking, and enabling the meta-analysis of AI-generated scientific work. The underlying technologies include template-mapping tools, submission-API integration, metadata provenance logging, and modules for ethical-disclosure verification.
4 Applications of AI Scientist Systems
AI Scientist systems are rapidly transitioning from conceptual frameworks to practical infrastructures for automated research. By autonomously generating hypotheses, designing experiments, analyzing data, and producing manuscripts, they are beginning to reshape the scientific landscape. To capture the field’s diversity, we categorize existing systems into two primary tiers: (1) General AI Scientist Systems, which pursue end-to-end, cross-domain scientific autonomy; and (2) Domain-Specific AI Scientist Systems, which are specialized for concrete scientific areas.
4.1 General AI Scientist Systems
General AI Scientist systems represent the most ambitious frontier in autonomous research, aiming to create domain-agnostic frameworks that computationally embody the scientific method itself. The goal is not merely to automate specific tasks, but to replicate the entire cognitive workflow of a human researcher: from identifying a promising research question to planning and executing experiments, interpreting the results, and communicating the findings. These systems serve a dual purpose: they are powerful tools for accelerating discovery, and they function as epistemological probes into the nature of machine intelligence and non-human scientific reasoning. The cornerstone architectures below define the state of the art in this pursuit.
  • $\bullet$
    The AI Scientist v1 [15] initiated one of the first truly autonomous, end-to-end research pipelines. Its modular architecture features distinct agents for planning, coding, analysis, and writing, all coordinated by a meta-scientist module. It demonstrated the ability to autonomously select research topics, generate executable code, and draft complete papers, successfully reproducing canonical machine learning studies.
  • $\bullet$
    The AI Scientist v2 [6] significantly advanced its predecessor by replacing a linear workflow with agentic tree-search planning. This allows the system to dynamically explore multiple research hypotheses in parallel, evaluating each for novelty and validity. The integration of a reflective feedback loop marks a critical step towards self-improving research agents.
  • $\bullet$
    AI-Researcher [46] is architected with a primary focus on transparency and verifiability. Its multi-agent system is underpinned by a provenance-tracking memory graph that records every intermediate artifact, from code to data logs. The framework is co-developed with its own benchmark to explicitly evaluate reproducibility and documentation quality.
  • $\bullet$
    Curie [44] concentrates on achieving rigorous experimental control within the "AI for AI research" paradigm. It employs causality-aware planning loops to automate the empirical testing of machine learning hypotheses, ensuring that each experimental choice is linked to explicit causal assumptions and bridging general reasoning with formal scientific standards.
4.2 Chemistry and Materials Science
The fields of chemistry and materials science have emerged as the earliest and most mature testbeds for AI Scientist systems. This maturity stems from a unique confluence of factors: highly structured representations for molecules, well-defined experimental procedures, and the increasing prevalence of robotic "self-driving laboratories" (SDLs). This environment is ideal for closing the loop between digital reasoning and real-world action to achieve de novo discovery. The systems below exemplify this evolution.
  • $\bullet$
    Coscientist [38] pioneered the integration of large language models with physical experimentation in chemistry. The system leverages a GPT-4 reasoning engine to autonomously plan reactions, write Python code to control robotic liquid handlers, and interpret spectroscopic feedback to verify outcomes and iteratively refine its hypotheses.
  • $\bullet$
    A-Lab [117] applies the principles of autonomous discovery to materials science. It integrates Bayesian optimization for efficient exploration of the parameter space with LLM-guided reasoning for experiment design. This synergy enables the system to autonomously synthesize and characterize thousands of novel inorganic materials per week.
  • $\bullet$
    Robotic AI Chemist [118] presents a fully automated, multi-agent robotic platform that performs end-to-end reaction design and on-demand physical execution. The system embodies the convergence of cognitive autonomy (literature-grounded reasoning) and physical autonomy (robotic manipulation), closing the loop in a real-world laboratory setting.
  • $\bullet$
    AutoLabs [31] introduces a multi-agent cognitive control architecture featuring "self-correction cycles." Specialized agents for planning, control, and safety auditing collaborate, enabling the system to detect experimental anomalies and automatically recalibrate laboratory instruments, thereby improving throughput and safety.
4.3 Biology and Biomedical Research
Biology and biomedical research present a domain of immense complexity, characterized by high stochasticity, intricate context-dependencies, and a vast, multimodal data landscape. The core challenge here is not just procedural automation, but the semantic interpretation of complex protocols and the inference of causal mechanisms from noisy, high-dimensional data. The ultimate ambition is to accelerate the discovery of disease pathways and identify novel drug targets. The following frameworks showcase key efforts in this area.
  • $\bullet$
    BioPlanner [39] established one of the earliest and most influential benchmarks for evaluating LLMs on the task of biological protocol design. It formalized the challenge of translating high-level research goals into complete, valid, and executable experimental workflows, providing a quantitative foundation for the field.
  • $\bullet$
    LLM4GRN [119] showcases the application of LLMs to a core challenge in systems biology: discovering causal gene regulatory networks. The system integrates LLM reasoning with external bioinformatics tools, demonstrating how AI can automate complex data analysis pipelines to infer biological mechanisms from experimental data.
  • $\bullet$
    Hierarchically Encapsulated Representation [90] addresses the complexity of biological protocols by introducing a hierarchical architecture for procedural design. This allows agents to reason about experiments at multiple levels of abstraction, from macro-level workflows down to micro-level parameter settings, enabling more robust and context-aware planning in self-driving labs.
4.4 Physics and Engineering
Physics and engineering represent a frontier where the goal for AI Scientists transcends data interpretation to encompass one of the ultimate scientific acts: the discovery of fundamental governing equations. The grand challenge in this domain is to perform abduction—to infer the underlying symbolic principles that govern a physical system from observational data. This requires a sophisticated synergy between numerical simulation, symbolic reasoning, and the real-time control of physical instrumentation. The systems below represent key advances in this pursuit.
  • $\bullet$
    Agentic Physics Experiments [94] demonstrated the groundbreaking deployment of AI agents for controlling multi-stage physics experiments at a large-scale particle accelerator facility. The system autonomously coordinated data acquisition and beamline configuration, using a closed feedback loop to optimize experimental parameters and significantly reduce calibration time.
  • $\bullet$
    SR-Scientist [19] tackles the classic scientific discovery task of symbolic regression. It employs an agentic AI workflow to autonomously discover fundamental scientific equations from observational data, a core competency that directly emulates a hallmark of human scientific intelligence from Galileo to Newton.
  • $\bullet$
    AI Feynman [120] is a foundational work in AI-driven physics discovery that introduced a paradigm of symbolic search guided by physical constraints (e.g., dimensional consistency). This approach successfully rediscovered several classical physical laws from raw data, inspiring the symbolic reasoning components in many modern AI Scientist systems.
  • $\bullet$
    Quantum-Agent-SDL [43] applies the principles of autonomous experimentation to the highly complex domain of quantum computing. The system uses a combination of reinforcement learning and LLM-based hypothesis refinement to perform self-optimized qubit calibration and error correction.
4.5 Meta-Science and Social Science
In this emerging and reflexive application area, the analytical lens of the AI Scientist is turned inward to study the structure, dynamics, and governance of the scientific enterprise itself. Instead of investigating the natural world, these systems analyze science as a complex adaptive system, mapping knowledge flows, identifying emerging paradigms, and assessing research reproducibility. This represents an epistemological turning point, where AI systems evolve from being executors of research to meta-researchers.
  • $\bullet$
    SciAgents [65] is a pioneering framework designed for the automated analysis of scientific knowledge graphs. It employs multi-agent collaboration and dynamic graph reasoning to traverse publication and author networks, enabling it to identify latent knowledge gaps, interdisciplinary connections, and predict future research trends.
  • $\bullet$
    AI for Social Science (AI4SS), as surveyed in [121], outline a roadmap for applying AI to social science. This body of research details how AI systems can be used for large-scale social modeling, simulating policy effects, and analyzing the diffusion of innovations and scientific collaboration patterns.
  • $\bullet$
    Ethical Governance Frameworks, as explored in a body of literature [109, 108, 110], are crucial for the responsible development of AI Scientists. This research focuses on the societal and ethical implications of AI-generated science, proposing frameworks for managing authorship, ensuring credit attribution, and maintaining accountability.
5 Open Problems and Future Directions
Despite the rapid progress chronicled in this survey, the journey toward a truly autonomous, general-purpose, and accountable AI Scientist is still in its nascent stages. Current systems, while impressive, exhibit limitations in robustness, generalizability, and trustworthiness. Based on our analysis, we identify four critical and interdependent frontiers for future research that must be addressed to advance the field from promising demonstrations to indispensable scientific partners.
From Reproducibility-by-Design to Verifiable Science.
The challenge of reproducibility is amplified in complex, multi-stage autonomous systems where minor variations can cascade into divergent outcomes. While benchmarks like BLADE [80] have introduced trace logging, a more fundamental paradigm of verifiable science is required. Future systems must be architected with reproducibility-by-design, incorporating three key elements: (1) environmental determinism, including containerized dependencies and hashed data artifacts; (2) fine-grained provenance, where every claim in a generated manuscript is cryptographically linked to the specific code, data, and model version that produced it; and (3) automated verification, where lightweight formal methods or AI-driven "auditor" agents check for logical consistency and claim-evidence alignment before publication, inspired by frameworks like SPOT [33].
Reasoning Under Uncertainty and Epistemic Humility.
Modern AI Scientists often produce outputs with a veneer of confidence, masking the underlying uncertainties, alternative hypotheses, or methodological trade-offs inherent in the research process. The next generation of systems must treat uncertainty as a first-class citizen. This requires moving beyond heuristic self-correction toward architectures that explicitly model and propagate uncertainty throughout the scientific workflow. Promising directions include the use of Bayesian deep learning, the ability to maintain and reason over multiple competing hypotheses in parallel (as seen in The AI Scientist v2 [6]), and the development of agents that demonstrate epistemic humility—knowing when to express low confidence, when to seek more data, and when to defer to human experts.
Cross-Domain Generalization through Modular and Composable Architectures.
While AI Scientists have shown remarkable success in highly structured domains like chemistry [38], they struggle to generalize to fields with less formalized procedures or rapidly evolving instrumentation. The current monolithic, end-to-end training paradigm is brittle. A more robust approach lies in developing modular and composable capabilities. This involves creating a standardized "toolkit" of expert modules—for causal inference, symbolic regression, data visualization, etc.—that can be dynamically orchestrated by a master planning agent. Such a capability-factorized architecture, where modules with explicit I/O contracts can be composed on the fly, would facilitate procedural transfer and allow an AI Scientist to tackle novel problems in unfamiliar domains by assembling known skills in new ways.
The Human-AI Collaborative Frontier and Ethical Governance.
As the frontier of autonomous capability expands, so does the importance of the human-AI interface. The most impactful scientific work in the near future will likely arise not from fully autonomous systems, but from deep, synergistic human-AI collaboration. The development of frameworks like Freephdlabor [36] points toward a future of interactive science automation, where the human researcher acts as a strategic director, guiding the AI’s exploration and validating its most creative leaps. This paradigm requires the development of novel, role-aware collaboration protocols and auditable interfaces. Concurrently, as AI-generated claims enter the scientific record, robust **ethical governance** becomes non-negotiable. Systems must embed machine-readable author contribution statements, transparent audit trails, and risk-gated execution for high-stakes research. As argued in the literature [109, 108], establishing community-wide standards for AI authorship and accountability is not just a technical challenge, but a societal necessity.
6 Conclusion
In this survey, we have charted the rapid and transformative evolution of the AI Scientist, tracing its progression from a fragmented landscape of specialized tools to a more coherent field of integrated, end-to-end research agents. By introducing a unified six-stage methodological framework and a three-phase historical narrative, we provided a durable conceptual scaffolding to systematically map and analyze the key systems, benchmarks, and applications that define this emerging paradigm. Our synthesis reveals a clear trajectory towards increasingly integrated, self-reflective, and autonomous systems. The current frontier is characterized by a dual thrust: the pursuit of greater scalability and scientific impact through systems like DeepScientist [34], and the simultaneous development of sophisticated human-AI collaborative frameworks like Freephdlabor [36]. While formidable challenges in generalizability and ethical governance remain, the ultimate promise of the AI Scientist lies not in replacing the human researcher, but in forging a new, symbiotic partnership. This collaboration is poised to augment human creativity with the scale, speed, and novel exploratory capabilities of intelligent machines, fundamentally redefining the nature of scientific inquiry and accelerating the pace of discovery for generations to come.

### v2

@@
 Through this analytical lens, we systematically map and analyze dozens of seminal works from 2022 to late 2025, revealing a clear three-phase evolutionary trajectory.
-Our analysis charts the field’s progression from an initial phase of Foundational Modules (2022–2023), focused on task-specific automation, through a period of Closed-Loop Integration (2024), to the current frontier of Scalability, Impact, and Collaboration (2025–present).
+Our analysis charts the field’s progression from an initial phase of Foundational Modules, focused on task-specific automation, through a period of Closed-Loop Integration, to the current frontier of Scalability, Impact, and Collaboration.
 By synthesizing these developments, this survey identifies key architectural patterns and highlights the dual research thrusts toward both greater machine autonomy and more sophisticated human-in-the-loop synergy.
@@
 This survey focuses on research between 2022 and 2025 that enables autonomous or near-autonomous execution of the scientific method.
-To ensure conceptual clarity and structural coherence, we adopt the radial organization illustrated in Figure 1, which conceptualizes the AI Scientist as a core engine driving a cyclical workflow that is instantiated across various scientific domains.
-At its heart lies the central concept of the Automatic AI Scientist System.
-Surrounding this core is an inner ring depicting the six fundamental methodological categories that constitute the scientific workflow: (1) Literature Review focuses on extracting, structuring, and reasoning over scientific corpora to establish prior knowledge foundations [27]; (2) Idea Generation addresses hypothesis formation and creative scientific reasoning [25, 26]; (3) Experimental Preparation encompasses data selection, variable definition, simulation setup, and statistical initialization [29, 30]; (4) Experimental Execution integrates protocol orchestration, robotic control, feedback-based iteration, and tool-use reasoning [12, 11, 31]; (5) Scientific Writing involves multi-modal evidence organization, figure and table generation, and result articulation [27, 32]; and (6) Paper Generation targets the synthesis of publishable, verifiable manuscripts that combine reasoning, visualization, and scholarly consistency [21, 33].
-As shown in Figure 1, encircling this workflow is an outer ring representing the application layers, showcasing where these methodologies are applied—from general-purpose frameworks to specialized domains such as Chemistry, Biology, Physics, and Meta-Science.
-This concentric structure provides a unified, navigable map for understanding how a core autonomous system executes a fundamental scientific process across a diverse and expanding landscape of applications.
-Figure 1: The concentric architecture of an AI Scientist System.
-The core represents the central Automatic AI Scientist System engine.
-This engine drives the six-stage methodological workflow, depicted in the inner ring, which forms a continuous, iterative cycle from Literature Review to Paper Generation.
-The outer ring illustrates the key application domains where this workflow is instantiated, from general-purpose frameworks to specialized fields like Chemistry, Biology, and Physics.
+To ensure conceptual clarity and structural coherence, we adopt the comprehensive architectural landscape illustrated in Figure 1 as the central organizing principle for our analysis.
+This structure is defined by two primary axes.
+The horizontal axis delineates the six sequential methodological stages of the scientific workflow: (1) Literature Review focuses on extracting, structuring, and reasoning over scientific corpora to establish prior knowledge foundations [27]; (2) Idea Generation addresses hypothesis formation and creative scientific reasoning [25, 26]; (3) Experimental Preparation encompasses data selection, variable definition, simulation setup, and statistical initialization [29, 30]; (4) Experimental Execution integrates protocol orchestration, robotic control, feedback-based iteration, and tool-use reasoning [12, 11, 31]; (5) Scientific Writing involves multi-modal evidence organization, figure and table generation, and result articulation [27, 32]; and (6) Paper Generation targets the synthesis of publishable, verifiable manuscripts that combine reasoning, visualization, and scholarly consistency [21, 33].
+The vertical axis defines four distinct layers of abstraction in a top-down hierarchy, flowing from high-level Application Domains, through concrete Scientific Tasks & Products and enabling System Architectures & Methods, down to the foundational Models & Capabilities.
+This 4x6 matrix provides a unified map for situating any given work by cross-referencing its methodological focus with its layer of contribution.
+Complementing this matrix, the figure also includes a timeline at the bottom that summarizes the field’s three-phase historical evolution, providing a chronological perspective on the progression from foundational modules to fully integrated systems.
+Figure 1: The Architectural Landscape of the Survey of AI Scientist for Automatic Research.
+The main 4x6 matrix maps the six methodological stages of the scientific workflow (horizontal axis) against four top-down layers of abstraction, from Applications to Models (vertical axis).
+The panel at the bottom illustrates the field’s three-phase historical evolution, categorizing representative works to provide a chronological perspective on the development of AI Scientist systems.
 1.1 Contributions